# Hi-Gen: Generative Retrieval For Large-Scale Personalized E-commerce Search

### v1

@@ Abstract.
Leveraging generative retrieval (GR) techniques to enhance search systems is an emerging methodology that has shown promising results in recent years. In GR, a text-to-text model maps string queries directly to relevant document identifiers (docIDs), so it dramatically simplifies the whole retrieval process. However, when applying most GR models in large-scale E-commerce for personalized item search, we have to face two key problems in encoding and decoding. (1) Existing docID generation methods ignore the encoding of efficiency information, which is critical in E-commerce. (2) The positional information is important in decoding docIDs, while prior studies have not adequately discriminated the significance of positional information or well exploited the inherent interrelation among these positions. To overcome these problems, we introduce an efficient Hierarchical encoding-decoding Generative retrieval method (Hi-Gen) for large-scale personalized E-commerce search systems. Specifically, we first design a representation learning model along with metric learning to learn discriminative feature representations of items to capture both semantic relevance and efficiency information. Then, we propose a category-guided hierarchical clustering scheme that makes full use of the semantic and efficiency information of items to facilitate docID generation. Finally, we design a position-aware loss to discriminate the importance of positions and mine the inherent interrelation between different tokens at the same position. This loss boosts the performance of the language model used in the decoding stage. Besides, we propose two variants of Hi-Gen (i.e.,Hi-Gen-I2I and Hi-Gen-Cluster) to support online real-time large-scale recall in the online serving process. Extensive experiments on both public and industry datasets demonstrate the effectiveness and efficiency of Hi-Gen. It gets 3.30 and 4.62 improvements over SOTA for Recall@1 on the public and industry datasets, respectively. Moreover, deploying Hi-Gen to a large-scale E-commerce platform leads to 6.89 and 1.42 improvements over the baseline model in the online AB-testing experiment in terms of recall and Gross Merchandise Value (GMV), respectively. In addition, Hi-Gen beats the basic Differentiable Search Index (DSI) model and BM25 in zero-shot learning scenarios, which proves its generalization capabilities.
1. Introduction
@@
With the rapid development of online E-commerce in recent years, personalized information retrieval technology has become increasingly important. For example, a user who requests a query on E-commerce websites like Amazon and AliExpress111AliExpress is one of the largest B2C cross-border platforms under Alibaba Group. can quickly retrieve a lot of relevant items from different seller stores. Information retrieval not only helps customers find their favorite items but also generates revenue for these business platforms.
@@
Existing retrieval techniques can be classified broadly into three categories: sparse retrieval (Lee et al., 2022; Dai and Callan, 2019; Zheng and Callan, 2015), dense retrieval (Karpukhin et al., 2020; Yu et al., 2021; Zeng et al., 2023; Wang et al., 2018; Zhang et al., 2020; Zhou et al., 2020), and generative retrieval (GR) (Sun et al., 2023). Sparse retrieval acquires documents by evaluating the relevance between query terms and document terms. Dense retrieval encodes queries and documents into high-dimensional vectors with a neural network and computes their similarity to retrieve relevant documents. In GR, the standardized pipelined retrieve-then-rank strategy for information retrieval is replaced by a simple sequence-to-sequence (Seq2Seq) strategy. The GR models in which generative Language Models (LMs) (Radford et al., 2019; Lewis et al., 2019; Liu et al., 2019; Yao et al., 2023) play a key role allow for end-to-end training and can serve as a differentiable sub-component in complex neural models for building modern intelligent E-commerce systems, which can enhance the customer experience. In light of this, GR has gained widespread attention in both academia and industry.
@@
Firstly, in the pre-processing stage, several term-based and vector-based strategies have been proposed to represent a document. For term-based approaches, a few representative words or tokens are selected to describe the semantic content of a document. In contrast, vector-based approaches present documents in dense vectors and model their similarities with inner product or cosine similarity, which effectively solve the problem of similar semantic meanings but different word representations faced by term-based methods. In addition to the semantic information, other side information, such as item conversion efficiency information, is also significant in e-commerce scenarios. Nevertheless, there are relatively few works that explicitly learn both the semantic and efficiency information in document222In the field of E-commerce search, we treat a product as a special document, so we sometimes use the word document or item to express a product when there is no ambiguity in this paper. representation. We call this problem the document representation problem.
@@
Secondly, document representations are mapped into their corresponding docIDs in the indexing stage. The existing docID generation methods can be roughly classified into three groups: 1) unstructured atomic identifiers, 2) naively structured string identifiers, and 3) semantically structured identifiers. The most naive way to represent documents is to assign each document an arbitrary (and possibly random) unique integer identifier. We refer to these representations as unstructured atomic identifiers following the work (Tay et al., 2022). Considering the limited expressive power of a single integer number, a more reasonable and clever way is to use a small group of arbitrary unique integers, words, or tokens to form docID. We refer to these as naively structured string identifiers. To capture some information about the semantics of its associated document, a hierarchical clustering process over document embeddings has been employed in (Tay et al., 2022) to introduce a decimal tree, which guides the generation of semantically structured identifiers. Structural representation has a wide range of potential applications because structural relationships exist in many real-world applications. However, the current docID generation methods of most GR models for large-scale E-commerce searches ignore encoding crucial efficiency information about items. We refer to this problem as the docID generation problem.
@@
Next, the retrieval process in GR is accomplished by decoding docIDs given an input query. When docIDs are formed as a series of organized tokens, decoding them is a structured prediction task where the position of each token is very important. A good structure prediction method should be able to capture dependencies between the output components and make joint predictions that respect these dependencies. It can lead to more accurate predictions compared to making independent predictions for each component. However, prior studies have not adequately discriminated the significance of positional information or well exploited the inherent interrelation among these positions (Lian et al., 2023). In a recent related work (Lian et al., 2023), Lian et al. began to consider the importance of different positions in docID by applying an additional hierarchical loss based on the document ID hierarchy at the decoding steps, but they ignored the inherent interrelation among different tokens at the same position. Here we use the term ”docID decoding problem” to describe this issue.
@@
Finally, directly using GR in online large-scale recall would encounter the service delay problem caused by the poor inference calculation performance. Most previous work in the field of GR only reported results on small datasets (Fan et al., 2023). Thus, there is a lack of real-world application testing results on large-scale industry applications such as large-scale E-commerce search and recommendation systems. In this paper, we call this issue the large-scale recall problem.
To overcome the above-mentioned problems, we propose an efficient Hierarchical encoding-decoding Generative retrieval method (Hi-Gen) for large-scale personalized E-commerce search systems. The main contributions of our work are:
@@
We introduce a novel algorithm for docID generation in GR. Our method first designs a representation learning model along with metric learning to learn a discriminative feature representation of a document for maintaining both semantic relevance and efficiency information. Subsequently, the category-guided hierarchical clustering scheme is leveraged to make full use of the semantic and efficiency information of documents to facilitate docID generation.
@@
The position-aware loss is proposed to enhance the performance of the language model used in the decoding stage. Our loss mechanism aims to discriminate the importance of positions and mine the semantic and efficiency differences among various tokens at the same position.
@@ 2. Related Work
Following the publication of Transformer Memory as a Differentiable Search Index(DSI) (Tay et al., 2022), a notable number of GR approaches have been proposed (Sun et al., 2023; Salemi et al., 2023; Zhou et al., 2022; Hua et al., 2023; Yu et al., 2022; Bevilacqua et al., 2022; Rajput et al., 2023; Wang et al., 2022; Zhao et al., 2022; Mehta et al., 2022; Zhuang et al., 2022; Zhou et al., 2023; Kishore et al., 2023; Raffel et al., 2020; Tang et al., 2023). The performance enhancement of GR can be attributed to three primary factors: firstly, through the refinement of the document representation; secondly, by enriching the docID encoding information; and lastly, by optimizing the employed language model.
@@
Document Representation. In (Rajput et al., 2023), a pre-trained content encoder was adopted to generate a semantic embedding for each item which has associated content features that capture useful semantic information. Lian et al. (Lian et al., 2023) utilized a transformer-based encoder-decoder network to jointly learn query embedding, document embedding, and personal embedding for learning the relationships among them. Wang et al. (Wang et al., 2022) introduced an additional query generation process to bind the information of document content through training the Seq2Seq model with generated queries and their corresponding document identifiers to make the identifiers aware of the document semantics. However, there are relatively few works that explicitly learn discriminative representation to encode both the semantic and efficiency information about the item.
@@
DocID Generation. Generally, a docID can be expressed as a single token (or atomic integers) or as a series of tokens, each of which could be a semantic numeric string or any arbitrary string. Tay et al. (Tay et al., 2022) investigated three kinds of docID generation strategies (i.e.,unstructured atomic identifiers, naively structured string identifiers, and semantically structured identifiers) and claimed that semantically structured identifiers lead to better indexing and retrieval capabilities for DSI. GenNet (Sun et al., 2023) learns to tokenize documents into semantic docIDs via a discrete auto-encoding approach. TIGER (Rajput et al., 2023) introduces a hierarchical quantizer (RQ-VAE) that generates semantic ID representations for items. SE-DSI (Tang et al., 2023) takes the generated pseudo query with explicit semantic meaning, called Elaborative Description, as the docID. Nevertheless, these methods solely focus on the semantic information of the documents but ignore the efficiency information. To capture both the semantic and efficiency information of items, we develop a discriminative representation process and then employ category-guided hierarchical clustering to generate docIDs.
@@
Language Model Optimization. To train the language model in GR, NCI (Wang et al., 2022) proposes a PAWA loss to tackle the significance variation problem when the same token appears in different places. PersonalTM (Lian et al., 2023) utilizes hierarchical loss to increase the impact of topper level clusters. Though they have acknowledged the specificity of the docID, the inherent semantic and efficiency information buried in different positions of the docID remains uncovered. Therefore, position-aware loss is suggested to distinguish the importance of different positions and discover the inherent interrelation between different tokens within the position. Since recent results suggest that some properties of large language models emerge only for very large model sizes, some existing work has tried to improve the LM by increasing model parameters or using better models (Tay et al., 2022).
3. Proposed Methodology
3.1. Overview
@@
Figure 1 depicts the overall architecture of Hi-Gen. Hi-Gen consists of offline and online stages. At the offline stage, a discriminative representation learning model is first employed to learn a fusion embedding for each item. Then the learned representation is encoded into a semantically and efficiently structured document identifier (docID) through a category-guided hierarchical clustering process. Subsequently, a pre-trained language model (LM) equipped with a position-aware loss is fine-tuned using the training data. The LM is fed with the user, query, and context-related features and learns to output the corresponding item docIDs. At the online stage, the fine-tuned LM model generates the most related top-K docIDs with a given query. Two variants of Hi-Gen (i.e.,Hi-Gen-I2I and Hi-Gen-Cluster) are designed to rapidly expand results based on top-K decoded docIDs to support large-scale recall. In what follows, we provide detailed descriptions of the proposed Hi-Gen.
3.2. Discriminative Representation Learning
@@
Using dense document representation to generate docID has been proven to be an effective method (Tay et al., 2022) in GR. However, the efficiency information about the item has never been taken into account in existing works. To eliminate this problem, we designed a discriminative representation learning model that captures both semantic relevance and efficiency information about the item. The model consists of two modules: 1) Atomic Embedding Learning Module, and 2) Metric Learning Module.
3.2.1. Atomic Embedding Learning Module
With user, query, item, and context features as input, three atomic embeddings (semantic, common, and efficient) are produced in this module. Figure 2 (a) illustrates the architecture of the atomic embedding learning module. For simplicity, we represent the output of the embedding layer for the user feature, query feature, and context feature in a batch as , , and , where and are embedding dimensions of each feature, and and are the sequence lengths for query and context, represents the batch size in training and inference stages.
To capture the underlying factors that explain variations in item data, we explicitly group the original features of an item into three parts: 1) item semantic feature: the semantic information about an item, such as its title; 2) item ID: unique identifier for an item; 3) item efficient feature: some efficiency-related statistical indicators, such as click and pay count within a week. After embedding and feature dimension reduction processing, these three parts of features are converted into three low-dimensional atomic features: , and , which are named as semantic embedding, common embedding, and efficient embedding, respectively. Notably, these three atomic item embeddings could have the same feature dimension (i.e.,). To capture both semantic and efficiency information of items, a multi-task two-tower embedding model (Yu et al., 2021) that consists of a user tower network and an item tower network is employed in which relevance prediction and click prediction tasks are deployed as downstream tasks as shown in Fig 2.
@@ 3.2.1. Atomic Embedding Learning Module
3.2.2. Metric Learning Module
Through the atomic embedding learning module, we acquire three atomic embeddings: , , and . However, since the two-tower network mainly focuses on modeling the similarity between item and user, the generated embeddings may be weak at distinguishing the semantic and efficiency information of similar items. As we know, metric learning can map the original features into some appropriate feature space where similar items are closer to each other, and dissimilar items are further apart (Hoffer and Ailon, 2015; Kaya and Bilge, 2019; Yang and Jin, 2006; Kulis et al., 2013). Thus, it is a suitable choice to introduce metric learning to refine the learned atomic embedding features.

### v2

@@ Abstract
Leveraging generative retrieval (GR) techniques to enhance search systems is an emerging methodology that has shown promising results in recent years. In GR, a text-to-text model maps string queries directly to relevant document identifiers (docIDs), dramatically simplifying the retrieval process. However, when applying most GR models in large-scale E-commerce for personalized item search, we must face two key problems in encoding and decoding. (1) Existing docID generation methods ignore the encoding of efficiency information, which is critical in E-commerce. (2) The positional information is important in decoding docIDs, while prior studies have not adequately discriminated the significance of positional information or well exploited the inherent interrelation among these positions. To overcome these problems, we introduce an efficient Hierarchical encoding-decoding Generative retrieval method (Hi-Gen) for large-scale personalized E-commerce search systems. Specifically, we first design a representation learning model using metric learning to learn discriminative feature representations of items to capture semantic relevance and efficiency information. Then, we propose a category-guided hierarchical clustering scheme that makes full use of the semantic and efficiency information of items to facilitate docID generation. Finally, we design a position-aware loss to discriminate the importance of positions and mine the inherent interrelation between different tokens at the same position. This loss boosts the performance of the language model used in the decoding stage. Besides, we propose two variants of Hi-Gen (i.e.,Hi-Gen-I2I and Hi-Gen-Cluster) to support online real-time large-scale recall in the online serving process. Extensive experiments on both public and industry datasets demonstrate the effectiveness and efficiency of Hi-Gen. It gets 3.30 and 4.62 improvements over SOTA for Recall@1 on the public and industry datasets, respectively. Moreover, deploying Hi-Gen to a large-scale E-commerce platform leads to 6.89 and 1.42 improvements over the baseline model in the online AB-testing experiment in terms of RecallNum and Gross Merchandise Value (GMV), respectively. In addition, Hi-Gen beats the basic Differentiable Search Index (DSI) model and BM25 in zero-shot learning scenarios, which proves its generalization capabilities.
Index Terms:
Search and Recommendation System, Information Retrieval, Generative Retrieval, Large Language ModelI Introduction
@@
With the rapid development of online E-commerce in recent years, personalized information retrieval technology has become increasingly important. For example, a user who requests a query on E-commerce websites like Amazon.com and AliExpress.com can quickly retrieve lots of relevant items from different seller stores. Information retrieval not only helps customers find their favorite items but also generates revenue for these business platforms.
@@
Roughly speaking, existing retrieval techniques can be classified broadly into three categories: sparse retrieval (SR) [1, 2], dense retrieval(DR) [3, 4, 5, 6, 7], and generative retrieval (GR) [8]. SR acquires documents by evaluating the relevance between query terms and document terms. DR encodes queries and documents into high-dimensional vectors with a neural network and computes their similarity to retrieve relevant documents. In GR, the standardized pipelined retrieve-then-rank strategy for information retrieval is replaced by a simple sequence-to-sequence (Seq2Seq) strategy. The GR models in which generative Language Models (LMs) [9, 10, 11, 12] play a key role allow for end-to-end training and can serve as a differentiable sub-component in complex neural models for building intelligent E-commerce systems to enhance the customer experience. In light of this, GR has gained widespread attention in recent years.
@@
Firstly, in the document representation stage, some term-based and vector-based strategies have been proposed to represent a document [8, 13]. For term-based approaches, a few representative words or tokens are selected to describe the semantic content of a document[13]. In contrast, vector-based approaches present documents in dense vectors and model their similarities with an inner product or cosine similarity, which effectively solves the problem of similar semantic meanings but different word representations faced by term-based methods[8]. In addition to the semantic information, other side information, such as item conversion efficiency information, is also significant in e-commerce scenarios. Efficiency information represents the efficiency-related indicators of the item, such as the Click-Through Rate (CTR) and Gross Merchandise Value (GMV). This information is very important for improving the efficiency of the personalized e-commerce search system. Nevertheless, there are relatively few works that explicitly learn both the semantic and efficiency information in document111In the field of E-commerce search, we treat a product as a special document, thus we sometimes use the word document or item to express a product when there is no ambiguity in this paper. representation. We call this problem the document representation problem.
@@
Secondly, document representations are mapped into their corresponding docIDs in the indexing stage. Existing docID generation methods can be roughly classified into three groups: 1) unstructured atomic identifiers, 2) naively structured string identifiers, and 3) semantically structured identifiers. The most naive way to represent documents is to assign each document an arbitrary (and possibly random) unique integer identifier, which is called unstructured atomic identifiers [8]. To enhance the representation ability, a small group of arbitrary unique integers, words, or tokens are used to form docID. We refer to these as naively structured string identifiers. To capture the semantic information about documents, a hierarchical clustering process over document embeddings has been employed in [8] to generate a decimal tree, which guides the generation of semantically structured identifiers. Structural representation has a wide range of potential applications because structural relationships exist in many real-world applications. However, the docID generation methods in most GR models ignore encoding crucial efficiency information about items in large-scale E-commerce search scenarios. We refer to this problem as the docID generation problem.
@@
Next, the retrieval process in GR is accomplished by decoding docIDs given a query. When docIDs are formed as a series of organized tokens, decoding them is a structured prediction task where the position of each token is very important. A good prediction method should capture dependencies between output components and make joint predictions that respect these dependencies, leading to more accurate predictions compared to making independent predictions for each component. However, prior studies have not adequately discriminated the significance of positional information or well exploited the inherent interrelation among these positions. Recently, Lian et al. [14] began to consider the importance of different positions in docID by applying an additional hierarchical loss at the decoding steps, but they ignored the inherent interrelation among different tokens at the same position. Here we use the term docID decoding problem to describe this issue.
@@
Finally, directly using GR in online large-scale recall applications would encounter the service delay problem caused by poor inference calculation performance. We call this issue the large-scale recall problem. Most previous research on GR only presented findings from small datasets [15]. There is a lack of real-world application testing results on large-scale applications such as industry E-commerce search systems.
To overcome the above-mentioned problems, we propose an efficient Hierarchical encoding-decoding Generative retrieval method (Hi-Gen). For document representation problem, Hi-Gen designs a representation learning model using metric learning to learn discriminative feature representations of items to capture both semantic relevance and efficiency information. To overcome docID generation problem, our method proposes a two-layer category-guided hierarchical clustering scheme, which facilitates the docID-related encoding and decoding process. As for docID decoding problem, a position-aware loss is proposed to boost the performance of the language model used in the decoding stage. Furthermore, we propose two variants of Hi-Gen (i.e.,Hi-Gen-I2I and Hi-Gen-Cluster) to solve the large-scale recall problem in real large-scale industry applications. Overall, the main contributions of our work are:
@@
We introduce a novel method for docID generation in GR. Our method first designs a representation learning model along with metric learning to learn a discriminative feature representation of a document for maintaining both semantic relevance and efficiency information. Subsequently, the category-guided hierarchical clustering scheme is adopted to make full use of the semantic and efficiency information of documents in docID generation.
@@
The position-aware loss is proposed to enhance the performance of the language model in the decoding stage. It aims to discriminate the importance of different positions and mine the semantic and efficiency differences among various tokens at the same position.
@@ II Related Work
Following the publication of Transformer Memory as a Differentiable Search Index(DSI) [8], a significant number of GR approaches have been proposed [13, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 13]. The performance improvement of General Relativity can be attributed to three primary factors: 1) refining the document representation; 2) enriching the docID encoding information; and 3) improving the language model used.
@@
Document Representaion. Wang et al. [22] introduced an additional query generation process for each document based on the content. Then they bound the information of document content through training the Seq2Seq model with generated queries and their corresponding document identifiers to make the identifiers aware of the document semantics. In [21], a pre-trained content encoder was adopted to generate a semantic embedding for each item which has associated content features that capture useful semantic information. Lian et al. [14] utilized a transformer-based encoder-decoder network to jointly learn query embedding, document embedding, and personal embedding for learning the relationship between the query and docID, document and docID, as well as the query and the personalized information. However, there are relatively few works that explicitly learn discriminative representation to encode both the semantic and efficiency information about the item.
@@
DocID Generation. Generally, a docID can be expressed as a single token (or atomic integers) or as a series of tokens, each of which could be a semantic numeric string or any arbitrary string. Tay et al. [8] investigated three kinds of docID generation strategies (i.e.,unstructured atomic identifiers, naively structured string identifiers, and semantically structured identifiers) and claimed that semantically structured identifiers lead to better indexing and retrieval capabilities for DSI. GenNet [29] learns to tokenize documents into semantic docIDs via a discrete auto-encoding approach. TIGER [21] introduces a hierarchical quantizer (RQ-VAE) that generates semantic ID representations for items. SE-DSI [13] takes the generated pseudo query with explicit semantic meaning, called Elaborative Description, as the docID. Nevertheless, these methods solely focus on the semantic information of the documents but ignore the efficiency information. To capture both the semantic and efficiency information of items, we develop a discriminative representation process and then employ category-guided hierarchical clustering to generate docIDs.
@@
Language Model Optimization. To train the language model in GR, NCI [22] proposes a PAWA loss to tackle the significance variation problem when the same token appears in different places. PersonalTM [14] utilizes hierarchical loss to increase the impact of topper level clusters. Though they have acknowledged the specificity of the docID, the inherent semantic and efficiency information buried in different positions of the docID remains uncovered. Therefore, position-aware loss is suggested to distinguish the importance of different positions and discover the inherent interrelation between different tokens within the position. Since recent results suggest that some properties of large language models emerge only for very large model sizes, some existing work has tried to improve the LM by increasing model parameters or using better models [8].
III Proposed Methodology
In this section, we introduce the proposed Hi-Gen model in detail. First, an overview of the model is illustrated. Subsequently, we clarify the procedures for discriminative representation learning. After that, the document identifier generation method is given. Then we introduce how to train the language model with position-aware loss. Finally, two variations of Hi-Gen are proposed to support real-time large-scale recall.
III-A Overview
@@
Figure 1 depicts the overall architecture of Hi-Gen. Hi-Gen consists of offline and online stages. At the offline stage, a discriminative representation learning model is first employed to learn a fusion embedding for each item. Then the learned representation is encoded into a semantically and efficiently structured document identifier (docID) through a category-guided hierarchical clustering process. Subsequently, a pre-trained language model (LM) equipped with a position-aware loss is fine-tuned using the training data. The LM is fed with the user, query, and context-related features and learns to output the corresponding item docIDs. At the online stage, the fine-tuned LM model generates the most related top-K docIDs with a given query. Two variants of Hi-Gen (i.e.,Hi-Gen-I2I and Hi-Gen-Cluster) are designed to rapidly expand results based on top-K decoded docIDs to support online large-scale recall. In what follows, we provide detailed descriptions of the proposed Hi-Gen.
III-B Discriminative Representation Learning
@@
Using dense document representation to generate docID has been proven to be an effective method [8] in GR. However, the efficiency information about the item has never been taken into account in existing works. To eliminate this problem, we designed a discriminative representation learning model that captures both semantic relevance and efficiency information about the item. The model consists of two modules: 1) Atomic Embedding Learning Module, and 2) Metric Learning Module.
III-B1 Atomic Embedding Learning Module
With user, query, item, and context features as input, three atomic embeddings (semantic, common, and efficient) are produced in this module, which represent semantic, universal and efficiency information of the item, respectively. Figure 2 (a) illustrates the architecture of the atomic embedding learning module. For simplicity, we represent the output of the embedding layer for the user feature, query feature, and context feature in a batch as , , and , where and are embedding dimensions of each feature, and and are the sequence lengths for query and context, represents the batch size in training and inference stages.
To capture the underlying factors that explain variations in item data, we explicitly group the original features of an item into three parts: 1) item semantic feature: the semantic information about an item, such as its title; 2) item ID: unique identifier for an item; 3) item efficient feature: some efficiency-related statistical indicators, such as click and pay count within a week. After embedding and feature dimension reduction processing, these three parts of features are converted into three low-dimensional atomic features: , and , which are named as semantic embedding, common embedding, and efficient embedding, respectively. Notably, these three atomic item embeddings could have the same feature dimension (i.e.,). To capture both semantic and efficiency information of items, a multi-task two-tower embedding model [4] that consists of a user tower network and an item tower network is employed in which relevance prediction and click prediction tasks are deployed as downstream tasks as shown in Fig 2.
@@ III-B1 Atomic Embedding Learning Module
III-B2 Metric Learning Module
Through the atomic embedding learning module, we acquire three atomic embeddings: , , and . However, since the two-tower network mainly focuses on modeling the similarity between item and user, the generated embeddings may be weak at distinguishing the semantic and efficiency information of similar items. As we know, metric learning can map the original features into some appropriate feature space where similar items are closer to each other, and dissimilar items are further apart [30, 31, 32, 33]. Thus, it is a suitable choice to introduce metric learning to refine the learned atomic embedding features.